# Geometric Signatures of Compositionality Across a Language Model’s Lifetime

### v2

@@ 5 Discussion
We have studied language model compositionality from a geometric and dynamic perspective. Using a carefully curated synthetic dataset, we found strong relationships between the compositionality of linguistic expressions and the dimensionality of their representations. On one hand, representational dimensionality is positively correlated to datasets’ formal, or combinatorial, complexity. On the other hand, grammatical sequences, whose semantics are composed via syntax, tend to exhibit a higher non-linear dimensionality, but a lower linear dimensionality, than agrammatical shuffled sequences. We showed that the positive relationship between compositionality and dimensionality is an inductive bias of the model, but that it is eventually shed in favor of learning a representation manifold that reflects meaningful semantic complexity in a phase transition. Results suggested differential coding of form and meaning in LM representations, where form complexity, estimated with gzip, is expressed linearly, and meaning complexity is expressed nonlinearly.
A central throughline in our results is that LMs compress representations to low-dimensional nonlinear manifolds, yet expand them to high-dimensional linear subspaces. This echoes independent results in computational neuroscience by Manley et al. (2024), who found that linear dimensionality scaled with number of neurons recorded in the mouse cortex, and De & Chaudhuri (2023), who found in neural networks that nonlinear, rather than linear, dimensionality better captured task semantics. The tendency for LMs to compress data to low-dimensional nonlinear manifolds, but, at the same time, expand them into high-dimensional linear subspaces, suggests a solution to the curse of dimensionality that also enjoys its blessings. High-dimensional representations classically engender overfitting and poor generalization (Hughes, 1968); but, these high-dimensional representations may lie on manifolds whose ID actually captures the latent sparsity of the data (De & Chaudhuri, 2023). At the same time, more dimensions implies more expressive orthogonality relations and linear decodability of categories (Cohen et al., 2020; Elmoznino & Bonner, 2023; Sorscher et al., 2022). Benefits of this dual patterning of intrinsic and effective dimensionality have been observed in biological and artificial intelligence (Jazayeri & Ostojic, 2021; Recanatesi et al., 2021; Haxby et al., 2011; Huth et al., 2012; De & Chaudhuri, 2023), where moreover, (linear) dimensional expansion and compression have implications for “lazy" and “rich" feature learning regimes, respectively (Flesch et al., 2022). While the present work is the first to show that this dual patterning in LMs broadly corresponds to a form-meaning dichotomy in representation learning, further work is needed to distentangle how nonlinear and linear features causally contribute to predictive coding.
Reproducibility statement
In Section 3.1, we have outlined the language models used in our experiments. The synthetic and naturalistic datasets employed to study compositionality and geometric feature complexity are introduced in Section 3.2 and further detailed in Appendix E. A comprehensive description of the measures used to assess representation geometric complexity is provided in Section 3.3, Appendix D, and Appendix C. Additionally, the benchmark tasks used to evaluate the Pythia checkpoints are summarized in Appendix F. Computing resources are described in Appendix A, and links to the assets used and their licenses are provided in Appendix B.
@@ Appendix A Computing resources
All experiments were run on a cluster with 12 nodes with 5 NVIDIA A30 GPUs and 48 CPUs each. Extracting LM representations took a few wall-clock hours per model-dataset computation. ID computation took approximately 0.5 hours per model-dataset computation. Taking parallelization into account, we estimate the overall wall-clock time taken by all experiments, including failed runs, preliminary experiments, etc., to be of about 10 days.
@@ Appendix B Assets
Appendix C Other Dimensionality Estimators
@@
In addition to TwoNN, we considered Levina & Bickel (2004)’s Maximum Likelihood Estimator (MLE), a similar, nonlinear measure of $I_{d}$. MLE has been used in prior works on representational geometry such as (Cai et al., 2021; Cheng et al., 2023; Pope et al., 2021), and similarly models the number of points in a neighborhood around a reference point $x$ to follow a Poisson point process. For details we refer to the original paper (Levina & Bickel, 2004). Like past work (Facco et al., 2017; Cheng et al., 2023), we found MLE and TwoNN to be highly correlated, producing results that were nearly identical: compare Figure 2 left to Figure G.3 left, and Figure G.1 top to Figure G.2 top).
@@
For our primary linear measure of dimensionality $d$, we computed PCA and took the number of components that explain 99% of the variance. In addition to PCA, we computed the Participation Ratio (PR), defined as $(\sum_{i}\lambda_{i})^{2}/(\sum_{i}\lambda_{i}^{2})$ (Gao et al., 2017). We found PR to give results that were incongruous with intuitions about linear dimensionality. In particular, it produced a lower dimensionality estimate than the nonlinear estimators we tested; see, e.g., Figure G.3, where the PR-$d$ for sane text is less than that of TwoNN. This contradicts the mathematical relationship that $I_{d}\leq d\leq D$. This may be because, empirically, PR-$d$ corresponded to explained variances of $60-80\%$, which are inadequate to describe the bounding linear subspace for the representation manifold. Therefore, while we report the mean PR-$d$ over model size in Figure G.3 and the dimensionality over layers in Figure G.2 for completeness, we do not attempt to interpret them.
Appendix D Intrinsic Dimension Details
ID estimation methods practically rely on a finite set of points and their nearest-neighbor structure in order to compute an estimated dimensionality value. The underlying geometric calculations assume that these are points sampled from a continuum, such as a lower-dimensional non-linear manifold. In our case, we actually have a discrete set of points so the notion of an underlying manifold is not strictly applicable. However, we can ask the question: if those points had been sampled from a manifold, what would the estimated ID be? Since the algorithms themselves only require a discrete set of points, they can be used to answer that question.
@@ Appendix E Controlled Grammar
The grammar is composed of sentences of the form
@@
Each category, colored and enclosed in brackets, is sampled from a vocabulary of 50 possible words, listed in the table below:
@@ Appendix F Benchmark tasks
Here we briefly summarize the benchmark tasks that we use to evaluate Pythia checkpoints as described in Section 4.3. In figure 3, we did not include WSC (Winogrande Schema Challenge) which was originally included in Biderman et al., as it has been proposed that WSC dataset performance on LMs might be corrupted by spurious biases in the dataset (Sakaguchi et al., 2021). Instead, we only presented the evaluation from WinoGrande task, which is inspired from original WSC task but adjusted to reduce the systematic bias (Sakaguchi et al., 2021).
@@ Appendix G Additional Results: Controlled Grammar
Figure G.1: Dimensionality over layers. TwoNN nonlinear $I_{d}$ (top) and PCA linear $d$ (bottom) over layers are shown for all sizes (left to right). Each color corresponds to a coupling length $k\in 1\cdots 4$. Solid curves denote sane sequences, and dotted curves denote shuffled sequences. For all models, lower $k$ results in higher $I_{d}$ and $d$ for both normal and shuffled settings. For all models, shuffling results in lower $I_{d}$ but higher $d$. Curves are averaged over 5 random seeds, shown with ±$\pm 1$ SD.
Figure G.2: Other dimensionality metrics over layers. MLE nonlinear $I_{d}$ (top) and PR linear $d$ (bottom) over layers are shown for all model sizes (left to right). Each color corresponds to a coupling length $k\in 1\cdots 4$. Solid curves denote sane sequences, and dotted curves denote shuffled sequences. For all models, lower $k$ results in higher $I_{d}$ for both normal and shuffled settings. For all models, shuffling results in lower $I_{d}$. The PR-$d$ produced nonsensical results, with linear dimensionality higher than nonlinear dimensionality. Curves are averaged over 5 random seeds, shown with ±$\pm 1$ SD.
@@
$\bm{\alpha}$ R p-value $\bm{\alpha}$ R p-value
sane 1 0.4598 0.9956 $2\times 10^{-6}$ 0.0023 0.6341 0.1261
sane 2 0.4268 0.9954 $3\times 10^{-6}$ 0.0011 0.5580 0.1930
sane 3 0.4014 0.9943 $5\times 10^{-6}$ 0.0009 0.6616 0.1056
sane 4 0.3569 0.9924 $1\times 10^{-5}$0.0003 -0.3523 0.4383
@@
Figure G.4: Layerwise feature complexity evolution over time, additional results. Nonlinear $I_{d}$ (top) and linear $d$ (bottom) over training is shown for sane (left) and shuffled (right) text, for the $1$-coupled setting. Each curve is one layer of the LM (yellow is later, purple is earlier). All settings in [TwoNN, PCA]$\times$[sane, shuffled] exhibit a phase transition in representational dimensionality at around checkpoint $10^{3}$, which corresponds to the sharp increase in task performance. In the nonlinear case (top row), the difference between layers’ $I_{d}$ is low at the end of training for shuffled text, and high for sane text. This suggests LM learns to perform meaningful and specialized processing over layers. The difference between layers’ $d$ (bottom row) at the end of training is, conversely, high for shuffled and lower for sane text. This is consistent with our interpretation of $d$ as capturing implied dataset size.
@@ Appendix H Additional Results: The Pile
$\bm{\alpha}$ R p-value $\bm{\alpha}$ R p-value
@@
Table H.1: Linear regression of average layerwise dimensionality on The Pile to hidden dimension, $D$. For dimensionality measures (PCA, TwoNN columns), the linear effect size $\alpha$ along with $R$-value and $p$-value are reported. PCA linear dimension shows a statistically significant linear relationship to $D$, with large effect size $\alpha=0.81$. TwoNN intrinsic dimension also shows a slightly weaker, but still highly significant, linear relationship to $D$. But, the effect size $\alpha$ is near-zero, showing that nonlinear $I_{d}$ is robust to changes in hidden dimension $D$.
@@
Figure H.2: ID phase transition in The Pile. Nonlinear $I_{d}$ (top) and linear $d$ (bottom) over training is shown for model sizes 410m, 1.4b, and 6.9b (left to right), for The Pile. Each curve is one layer of the LM (yellow is later, purple is earlier). Representations of The Pile exhibit a phase transition in both $I_{d}$ and $d$ at slightly before checkpoint $10^{3}$, where $t=10^{3}$ corresponds to a dip and redistribution of layerwise dimensionality, and also a sharp increase in task performance in Figure 3.
Appendix I Additional Results: Per-Layer Correlation with Kolmogorov Complexity

### v3

@@ 6 Discussion
On extensive experiments spanning degree of compositionality, sequence length, model size and training time, we found strong links between the compositionality of linguistic inputs and the complexity of their representations. We asked what aspects of feature complexity encode inputs’ (1) surface variation and (2) compositional structure, finding that linear $d$ tracks (1) and nonlinear $I_{d}$ tracks (2). We showed feature complexity to encode inputs’ superficial complexity as an inductive bias of the architecture; but that, over training, the bias is shed to reflect meaningful linguistic compositionality. This change marks an onset in linguistic competence, a cue for compositional understanding.
A central throughline in our results is that LMs represent their inputs on low-dimensional nonlinear manifolds, yet, at the same time, in high-dimensional linear subspaces. This tendency suggests a solution to the curse of dimensionality that also enjoys its blessings. High-dimensional representations classically engender overfitting and poor generalization (Hughes, 1968); but, they may form low-dimensional manifolds that capture the task’s latent, in our case, compositional, structure (De and Chaudhuri, 2023). At the same time, higher linear dimensionality has been shown to improve generalization thanks to its expressivity (Cohen et al., 2020; Elmoznino and Bonner, 2023; Sorscher et al., 2022). Intriguingly, the dual patterning of intrinsic and effective dimension, where latent structure is captured by low-$I_{d}$ manifolds that live in high-$d$ linear spaces, has been observed in the neuroscience literature for both biological and artificial systems (Jazayeri and Ostojic, 2021; Recanatesi et al., 2021; Haxby et al., 2011; Huth et al., 2012; De and Chaudhuri, 2023; Manley et al., 2024). That LMs also conform to this picture suggests that representation spaces that are both highly expressive ($d$) yet geometrically constrained ($I_{d}$) may be desirable in intelligent systems that recover compositional structure from data.
Limitations
  • While the present work is the first to show that nonlinear and linear representational dimension in LMs correspond to a form-meaning dichotomy, further work is needed to disentangle how nonlinear and linear features causally contribute to predictive coding. This would require careful probing and feature attribution.
  • Due to computational constraints, the present work is limited to several syntactic structures. Future work can apply this framework to investigate other structures like recursive embedding.
  • Also due to computational constraints, the present work is limited to model sizes of up to 8B parameters. We encourage further, better-resourced research to apply our framework to larger models.
  • Dimensionality measures tell us how complex the features are, but not what they are. How to isolate and interpret nonlinear features remains an open problem in the literature.
@@ Appendix A Computing resources
All experiments were run on a cluster with 12 nodes with 5 NVIDIA A30 GPUs and 48 CPUs each. Extracting LM representations took a few wall-clock hours per model-dataset computation. $I_{d}$ computation took approximately 0.5 hours per model-dataset computation. Taking parallelization into account, we estimate the overall wall-clock time taken by all experiments, including failed runs, preliminary experiments, etc., to be of about 10 days.
@@ Appendix B Assets
Appendix C Intrinsic Dimension Details
$I_{d}$ estimation methods practically rely on a finite set of points and their nearest-neighbor structure in order to compute an estimated dimensionality value. The underlying geometric calculations assume that these are points sampled from a continuum, such as a lower-dimensional non-linear manifold. In our case, we actually have a discrete set of points so the notion of an underlying manifold is not strictly applicable. However, we can ask the question: if those points had been sampled from a manifold, what would the estimated $I_{d}$ be? Since the algorithms themselves only require a discrete set of points, they can be used to answer that question.
Appendix D Other Dimensionality Estimators
@@
In addition to TwoNN, we considered Levina and Bickel (2004)’s Maximum Likelihood Estimator (MLE), a similar, nonlinear measure of $I_{d}$. MLE has been used in prior works on representational geometry such as (Cai et al., 2021; Cheng et al., 2023; Pope et al., 2021), and similarly models the number of points in a neighborhood around a reference point $x$ to follow a Poisson point process. For details we refer to the original paper (Levina and Bickel, 2004). Like past work (Facco et al., 2017; Cheng et al., 2023), we found MLE and TwoNN to be highly correlated, producing results that were nearly identical: compare Figure 1 left to Figure G.3 left, and Figure G.1 top to Figure G.4 top.
@@
For our primary linear measure of dimensionality $d$, we computed PCA and took the number of components that explain 99% of the variance. In addition to PCA, we computed the Participation Ratio (PR), defined as $(\sum_{i}\lambda_{i})^{2}/(\sum_{i}\lambda_{i}^{2})$ (Gao et al., 2017). We found PR to give results that were incongruous with intuitions about linear dimensionality. In particular, it produced a lower dimensionality estimate than the nonlinear estimators we tested; see, e.g., Figure G.3, where the PR-$d$ for coherent text is less than that of TwoNN. This contradicts the mathematical relationship that $I_{d}\leq d\leq D$. This may be because, empirically, PR-$d$ corresponded to explained variances of $60-80\%$, which are inadequate to describe the bounding linear subspace for the representation manifold. Therefore, while we report the mean PR-$d$ over model size in Figure G.3 and the dimensionality over layers in Figure G.4 for completeness, we do not attempt to interpret them.
@@ Appendix E Controlled Grammar
Figure E.1: Dataset structure and distributional properties. Top: The structure of the stimulus dataset. The top row shows the ordering of word categories, such as quality.ADJ or animal.N; below it, the vocabulary for each category, including words like “strong" (quality.ADJ) and “lion" (animal.N), respectively. When controlling the degree of dataset compositionality, contiguous word positions are coupled. For instance, when $k=3$, the first vocabulary indices for quality1.ADJ, nationality1.ADJ, and job1.N are tied together, such that “strong Thai doctor" or “calm French teacher" can be sampled, but “strong French doctor" cannot. Left: Examples of generated prompts for the normal, $k=3$, and shuffled settings. Right: When controlling the compositionality across $k=1\cdots 4$, word unigram frequencies are preserved in the resulting datasets, shown in the distributions looking identical.
Here, we describe the sampling procedure of the controlled grammar in more detail.
E.1 Sampling Procedure
Let there be $l$ variable words in the sequence (these are the colored words in the grammar in Section 3.1). Each position $i=1\cdots l$ is associated with a vocabulary $\mathcal{V}^{(i)}$ which contains vocabulary items $v_{j}^{(i)}$, $j=1\cdots|\mathcal{V}^{(i)}|$. We set $|\mathcal{V}^{(i)}|=50$ (line 153) for all $i$. All words in $\mathcal{V}^{(i)}$ are listed in the below Table.
We state the sampling procedure, illustrated in Figure E.1, as follows.
  • $\mathbf{k=1}$ For each position $i$ in the sentence, sample the word $w_{i}\sim\text{Unif}(\mathcal{V}^{(i)})$.
  • $\mathbf{k=2}$ Couples the vocabularies over bigrams in the sentence. Let $\circ$ be a concatenation operator. We construct a new coupled vocabulary $\mathcal{V}^{(i,i+1)}$ consisting of bigrams $v_{j}^{(i,i+1)}\triangleq v_{j}^{(i)}\circ v_{j}^{(i+1)}$, $j=1\cdots|\mathcal{V}^{(i)}|$. Then, for $i=1,3,5,\cdots l-1$, sample bigrams $w_{i}\circ w_{i+1}\sim\text{Unif}({\mathcal{V}^{(i,i+1)}})$.
  • $\mathbf{k=3}$ Couples trigrams in the sentence.
    Similarly, a new vocabulary over trigrams is constructed: $\mathcal{V}^{(i,i+1,i+2)}$ consists of trigrams $v_{j}^{(i,i+1,i+2)}:=v_{j}^{(i)}\circ v_{j}^{(i+1)}\circ v_{j}^{(i+2)}$. For $i=1,4,7,\cdots l-2$, sample trigrams $w_{i}w_{i+1}w_{i+2}\sim\text{Unif}({\mathcal{V}^{(i,i+1,i+2)}})$
and so on.
We design 5 different grammars of varying lengths $l\in\{5,8,11,15,17\}$ words.. The $17$-word grammar is the one used for all controlled grammar experiments except the "Varying Sequence Length" experiments (Appendix J). The structures of the grammars can be found below.
E.2 Length: 5 words
The [job1.N] [action1.V] the [animal.N].
E.3 Length: 8 words
The [nationality1.ADJ] [job1.N] [action1.V] the [color.ADJ] [texture.ADJ] [animal.N]
E.4 Length: 11 words
The [size2.ADJ] [quality1.ADJ] [nationality1.ADJ] [job1.N] [action1.V] the [size1.ADJ] [color.ADJ] [texture.ADJ] [animal.N]
E.5 Length: 15 words
The [quality1.ADJ] [nationality1.ADJ] [job1.N] [action1.V] the [size1.ADJ] [color.ADJ] [texture.ADJ] [animal.N] then [action2.V] the [size2.ADJ][job2.N].
E.6 Length: 17 words
@@
Each category, colored and enclosed in brackets, is sampled from a vocabulary of 50 possible words, listed in the table on the following page:
@@ Appendix F Benchmark tasks
Here we briefly summarize the benchmark tasks that we use to evaluate Pythia checkpoints as described in Section 4.3. In figure Figure 2, we $I_{d}$ not include WSC (Winogrande Schema Challenge) which was originally included in Biderman et al., as it has been proposed that WSC dataset performance on LMs might be corrupted by spurious biases in the dataset (Sakaguchi et al., 2021). Instead, we only presented the evaluation from WinoGrande task, which is inspired from original WSC task but adjusted to reduce the systematic bias (Sakaguchi et al., 2021).
@@ Appendix G Additional Results: Controlled Grammar
Here, we present extended results on the length-$17$ (longest) controlled grammar, for
  1. 1.
    All Pythia sizes, see Figure G.1, and Llama-3-8B and Mistral-7B, see Figure G.2.
  2. 2.
    Other dimensionality measures MLE (Levina and Bickel, 2004) and Participation Ratio (PR) (Gao et al., 2017), shown for Pythia models, see Figures G.3 and G.4. Results are highly consistent between TwoNN and MLE, but PR produced nonsensical values $d_{PR}<I_{d}$ that violated the theoretical relation $I_{d}\leq d\leq D$.
  3. 3.
    Linear fits $I_{d}\sim D$, $d\sim D$ for settings $k\in\{1\cdots 4\}\times\{\text{coherent},\text{shuffled}\}$, see Table G.1.
  4. 4.
    Training dynamics of feature complexity (TwoNN, PCA) for three Pythia sizes 410m, 1.4b, and 6.9b, on both coherent $k=1$ and shuffled text, see Figure G.5.
Figure G.1: Dimensionality over layers. TwoNN nonlinear $I_{d}$ (top) and PCA linear $d$ (bottom) over layers are shown for all sizes (left to right). Each color corresponds to a coupling length $k\in 1\cdots 4$. solid curves denote coherent sequences, and dotted curves denote shuffled sequences. For all models, lower $k$ results in higher $I_{d}$ and $d$ for both normal and shuffled settings. For all models, shuffling results in lower $I_{d}$ but higher $d$. Curves are averaged over 5 random seeds, shown with ±$\pm 1$ SD.
Figure G.2: Dimensionality over layers for Llama-3-8B and Mistral-7B. TwoNN nonlinear $I_{d}$ (top) and PCA linear $d$ (bottom) over layers are shown for all sizes (left to right). Each color corresponds to a coupling length $k\in 1\cdots 4$. solid curves denote coherent sequences, and dotted curves denote shuffled sequences. For all models, lower $k$ results in higher $I_{d}$ and $d$ for both normal and shuffled settings. For all models, shuffling results in lower $I_{d}$ but higher $d$. Curves are averaged over 5 random seeds, shown with ±$\pm 1$ SD. Results mirror those of the Pythia models.
Figure G.4: Other dimensionality metrics over layers. MLE nonlinear $I_{d}$ (top) and PR linear $d$ (bottom) over layers are shown for all model sizes (left to right). Each color corresponds to a coupling length $k\in 1\cdots 4$. solid curves denote coherent sequences, and dotted curves denote shuffled sequences. For all models, lower $k$ results in higher $I_{d}$ for both normal and shuffled settings. For all models, shuffling results in lower $I_{d}$. The PR-$d$ produced nonsensical results, with linear dimensionality higher than nonlinear dimensionality. Curves are averaged over 5 random seeds, shown with ±$\pm 1$ SD.
@@
$\boldsymbol{\alpha}$ R p-value $\boldsymbol{\alpha}$ R p-value
coherent 1 0.4598 0.9956 $2\times 10^{-6}$ 0.0023 0.6341 0.1261
coherent 2 0.4268 0.9954 $3\times 10^{-6}$ 0.0011 0.5580 0.1930
coherent 3 0.4014 0.9943 $5\times 10^{-6}$ 0.0009 0.6616 0.1056
coherent 4 0.3569 0.9924 $1\times 10^{-5}$0.0003 -0.3523 0.4383
@@
Figure G.5: Layerwise feature complexity evolution over time, additional results. Nonlinear $I_{d}$ (top) and linear $d$ (bottom) over training is shown for coherent (left) and shuffled (right) text, for the $1$-coupled setting. Each curve is one layer of the LM (yellow is later, purple is earlier). All settings in [TwoNN, PCA]$\times$[coherent, shuffled] exhibit a phase transition in representational dimensionality at around checkpoint $10^{3}$, which corresponds to the sharp increase in task performance. In the nonlinear case (top row), the difference between layers’ $I_{d}$ is low at the end of training for shuffled text, and high for coherent text. This suggests LM learns to perform meaningful and specialized processing over layers. The difference between layers’ $d$ (bottom row) at the end of training is, conversely, high for shuffled and lower for coherent text. This is consistent with our interpretation of $d$ as capturing implied dataset size.
@@ Appendix H Additional Results: The Pile
Similar to for the controlled grammar (Section 4.1), representational $I_{d}$ and $d$ scale differently with model ambient dimension $D$, where we found for the controlled grammar that $d\propto D$ while $I_{d}$ is stable in $D$. Here, Figure H.1 and Table H.1 similarly show that $d\propto D$ with a highly significant linear fit; a large effect $\alpha=0.81$ indicates the LM fills the ambient space such that $d\approx D$. For The Pile, $I_{d}\propto D$ ($R=0.95$, $p<0.001$) as well, but the tiny effect $\alpha$=$0.002$ shows $I_{d}$ is stable as $D$ grows, see Figure H.1 (left).
@@
$\boldsymbol{\alpha}$ R p-value $\boldsymbol{\alpha}$ R p-value
@@
Table H.1: Linear regression of Pythia’s average layerwise dimensionality on The Pile to hidden dimension, $D$. For dimensionality measures (PCA, TwoNN columns), the linear effect size $\alpha$ along with $R$-value and $p$-value are reported. PCA linear dimension shows a statistically significant linear relationship to $D$, with large effect size $\alpha=0.81$. TwoNN intrinsic dimension also shows a slightly weaker, but still highly significant, linear relationship to $D$. But, the effect size $\alpha$ is near-zero, showing that nonlinear $I_{d}$ is robust to changes in hidden dimension $D$.
@@
Figure H.2: $I_{d}$ phase transition in The Pile. Nonlinear $I_{d}$ (top) and linear $d$ (bottom) over training is shown for model sizes 410m, 1.4b, and 6.9b (left to right), for The Pile. Each curve is one layer of the LM (yellow is later, purple is earlier). Representations of The Pile exhibit a phase transition in both $I_{d}$ and $d$ at slightly before checkpoint $10^{3}$, where $t=10^{3}$ corresponds to a dip and redistribution of layerwise dimensionality, and also a sharp increase in task performance in Figure 2.
Appendix I Additional Results: Correlation with Kolmogorov Complexity
I.1 Correlation between Kolmogorov Complexity and Feature Complexity is Robust to Sequence Length
In Table 1 we showed that, for each model, and on a single dataset ($k=1$, $l=17$), linear effective $d$ highly correlates to the estimated superficial complexity (KC) using gzip. Table I.1 shows that this trend is robust to both model family, model size, and sequence length; average layerwise $d$ is almost perfectly monotonic in the KC of the dataset, seen by high Spearman correlation. In contrast, for none of the sequence lengths is average layerwise $I_{d}$ monotonic in KC, except for the smallest Pythia model (14m).
I.2 Kolmogorov Complexity vs. Average-Layer Feature Complexity Across Datasets
Figure I.1 shows the global correlation between feature complexity ($I_{d}$ and $d$) and KC, estimated with gzip. While both nonlinear (top row) and linear (bottom row) dimensionality are positively Spearman-correlated to gzip, there are clear differences:
  1. 1.
    Linear $d$ increases in the shuffled setting from the coherent setting; nonlinear $I_{d}$ decreases.
  2. 2.
    Linear $d$ is very highly correlated to the estimated KC, $\rho\approx 0.9$ in all cases, while nonlinear $d$ is more weakly correlated, $\rho\in[0.5,0.6]$.
These observations support the hypothesis that linear effective $d$ encodes KC, while the intrinsic dimension $I_{d}$ encodes sequence-level semantic complexity.